OpenStudy (anonymous):

are vectors [1, 0 ,1 ], [1, 1, 1] and [1, 0,-1] coplanar

OpenStudy (amistre64):

2 of them are in the same plane, but that middle one seems to be sticking out of the box

OpenStudy (amistre64):

... but then I could be sooo wrong too ;)

OpenStudy (amistre64):

i still think the 1st and 3rd are planar but the 2nd one sticks off to the side

OpenStudy (zarkon):

what is the span of these 3 vectors

OpenStudy (anonymous):

A=Bi+CJ [1,1,1]=i[1,0,1]+j[1,0,-1] 1=i+j 1=0j+0j -----> 1=0 nope 1=1i-1j that's what I think, they are not coplanar

OpenStudy (zarkon):

they are not coplanar.